# FireANTs: Adaptive Riemannian Optimization for Multi-Scale Diffeomorphic Registration

### v1

@@ 4 Methods
$\varphi_{k}=\begin{cases}\text{Id}&\text{if }k=1\\ \text{Upsample}(\varphi_{k-1})&\text{otherwise}\end{cases}$
@@
Unlike upsampling images, upsampling warp fields and their corresponding optimizer state variables requires careful consideration of the interpolation strategy. Bicubic interpolation is a commonly used strategy for upsampling images to preserve smoothness and avoid aliasing. However, bicubic interpolation of the warp field can lead to overshooting, leading to introducing singularities in the upsampled displacement field when there existed none in the original displacement field. In contrast, bilinear or trilinear interpolation does not lead to overshooting, and therefore diffeomorphism of the upsampled displacement is guaranteed, if the original displacement is diffeomorphic.
@@
We demonstrate this using a simple 2D warp field in  Fig. S.3(b). On the left, we consider a warp field created by nonlinear shear forces. This warp field does not contain any tears or folds - and is diffeomorphic. We upsample this warp field using bicubic interpolation (top) and bilinear interpolation (bottom). We also plot a heatmap of the negative of the determinant of the Jacobian of the upsampled warp, with a contour representing the zero level set. Qualitatively, bicubic interpolation introduces noticable folds in the warping field, leading to non-diffeomorphisms in the upsampled warp field. The heatmap shows a significant portion of the upsampled warp field has a negative determinant, indicating non-invertibility. On the other hand, bilinear interpolation looks blocky but preserves diffeomorphism everywhere, as also quantitatively verified by the absence of a zero level set in the heatmap.
Modular software implementation to enable effective experimentation
@@
Our software is implemented completely using default primitives in PyTorch. All code and example scripts is available at https://github.com/rohitrango/fireants.
4.1 Experiment Setup
Klein et al. brain mapping challenge  5
Brain mapping requires a common coordinate reference frame to consistently and accurately communicate the spatial relationships within the data. Automatically determining anatomical correspondence is almost universally done by registering brains to one another or to a template. Klein et al.evaluate a suite of fully automated nonlinear deformation algorithms applied to human brain image registration. A natural way to evaluate whether two images are in a common coordinate frame is to evaluate the accuracy of overlap of gross morphological structures (gryi, sulci, subcortical regions for example). The evaluation considers a total of four T1-weighted brain datasets with different whole-brain labelling protocols, eight different evaluation measures and three independent analysis methods. The paper evaluates 14 nonlinear registration algorithms with different parameterizations and assupmtions about the deformation field, and different regularizations.
Brain image data and their corresponding labels for 80 normal subjects were acquired from four different datasets. The LPBA40 dataset contains 40 brain images and their labels to construct the LONI Probabilistic Brain Atlas (LPBA40). All volumes were skull-stripped, and aligned to the MNI305 atlas  88 using rigid-body transformation to correct for head tilt. For all these subjects, 56 structures were manually labelled and bias-corrected using the BrainSuite software. The IBSR18 dataset contains brain images acquired at different laboraties through the Internet Brain Segmentation Repository. The T1-weighted images were rotated to be in Talairach alignment and bias-corrected. Manual labelling is performed resulting in 84 labeled regions. For the CUMC12 dataset, 12 subjects were scanned at Columbia University Medical Center on a 1.5T GE scanner. Images were resliced, rotated, segmented and manually labeled, leading to 128 labeled regions. Finally, the MGH10 dataset contains 10 subjects who were scanned at the MGH/MIT/HMS Athinoula A. Martinos Center using a 3T Siemens scanner. The data is bias-corrected, affine-registered to the MNI152 template, and segmented. Finally the images were manually labeled, leading to 74 labeled regions. All datasets have a volume of $256\times 256\times\{128,124\}$ voxels with varying amounts of anisotropic voxel spacing, ranging from $0.84\times 0.84\times 1.5$mm to $1\times 1\times 1.33$mm.
ANTs was one of the top performing methods for this challenge, performing well robustly across all four datasets. The method considers measures of volume and surface overlap, volume similarity, and distance measures to evaluate the alignment of anatomical regions. Given a source label map ${S}_{r}$ and target label map ${T}_{r}$ and a cardinality operator |.||.|| . |, we consider the following overlap measures. The first measure ‘target overlap’, defined as the overlap between the source and target divided by the target.
$TO_{r}=\frac{|{S}_{r}\cap{T}_{r}|}{|{T}_{r}|}$ (4)
Target overlap is a measure of sensitivity, and the original evaluation 5 considers the aggregate total overlap as follows
$TO_{Klein}=\frac{\sum_{r}|{S}_{r}\cap{T}_{r}|}{\sum_{r}|{T}_{r}|}$ (5)
$TO=\frac{1}{N_{r}}\sum_{r}TO_{r}$ (6)
$MO_{r}$ $=2\frac{|{S}_{r}\cap{T}_{r}|}{|{S}_{r}|+|{T}_{r}|}$ (7)
$MO_{Klein}$ $=2\frac{\sum_{r}|{S}_{r}\cap{T}_{r}|}{\sum_{r}(|{S}_{r}|+|{T}_{r}|)}$ (8)
$MO$ $=\frac{1}{N_{r}}\sum_{r}MO_{r}$ (9)
Klein et al.5 also propose a ‘Union Overlap’ metric which is a monotonic function of the Dice score. Therefore, we do not use this in our evaluation. To complement the above agreement measures, we also compute false negatives (FN), false positives (FP), and volume similarity (VS) coefficient for anatomical region $r$:
Similar to the overlap metrics, we compute the aggregates as in the original evaluation denoted by $FN_{Klein},FP_{Klein},VS_{Klein}$ and average over regions denoted simply by FN, FP, VS. This leads to a total of 10 aggregate metrics that we use to compare our method with 4 baselines - ANTs, Demons, VoxelMorph and SynthMorph.
EMPIRE10 challenge  10
Alignment of thoracic CT images, especially the lung and its internal structures is a challenging task, owing to the highly deformable nature of the lungs. Pulmonary registration is clinically useful, for example registering temporally distinct breathhold scans make visual comparison of these scans easier and less error prone. Registering inspiration and expiration scans can also be used to model or understand the biomechanics of lung expansion. Registration of temporally spaced breathhold scans can help in tracking disease progression, or registration between inspiration and expiration scans can enable improved monitoring of airflow and pulmonary function. Murphy et al.propose the Evaluation of Methods for Pulmonary Image REgistration 2010 (EMPIRE10) challenge to provide a platform for a comprehensive evaluation and fair comparison of registration algorithms for the task of CT lung registration. The dataset consists of 30 scan pairs including inspiration-expiration scans, breathhold scans over time, scans from 4D data, ovine data, contrast-noncontrast, and artificially warped scan pairs. The ovine data was acquired where breathing was controlled, and metallic markers were surgically implanted to provide landmark annotations, followed by a hole-filling algorithm to disguise the markers so that registration algorithms cannot use this artificial information. Artificially warped scan pairs also provide ground truth correspondences for landmarks and lung boundaries. The challenge provides a broad range of data complexity, voxel sizes and image acquisition differences. In this challenge, only intrapatient registration is considered, and lungs and lung fissures were segmented using an automated method, and altered manually wherever necessary. The challenge only provides scan pairs and binary lung masks. All the other data (fissures and landmarks) are withheld for evaluation. All the scan pairs have varying spatial and physical resolutions, are acquired over a varying set of imaging configurations. This calls for a registration algorithm that is agnostic to any assumptions about anisotropy of image resolution, both physical and voxel. We use the evaluation provided by the challenge, and compare the fissure alignment, landmark alignment, and singularity of registration. More details about the evaluation can be found in  10. We compare our method with ANTs which performs direct gradient descent updates and DARTEL which optimizes a stationary velocity field using the metrics reported in the evaluation server.
RnR ExM mouse dataset  64
[…]
Appendix A Extended Data
Refer to captionRefer to captionRefer to captionRefer to caption
@@ Appendix A Extended Data
Figure S.1: Regionwise target overlap on the brain MRI datasets: We further evaluate regionwise overlap scores by sampling 15 regions from each dataset, and comparing their distribution using our method and ANTs. Our method has a much higher median score, and better interquartile ranges across regions, demonstrating both accuracy and robustness.
@@
Figure S.2: Comparison of our method with ANTs on 4 MRI brain datasets: Registration quality is validated by measuring volume overlap of label maps between the fixed and warped label maps. (a): For anatomical region $r$, warped (binary) label map ${S}_{r}$ and fixed label map ${T}_{r}$, target and mean overlap are defined as $|{S}_{r}\cap{T}_{r}|/|{T}_{r}|$ and $2|{S}_{r}\cap{T}_{r}|/(|{S}_{r}|+|{T}_{r}|)$. We define the aggregate target overlap over all anatomical regions as $\sum_{r}(|{S}_{r}\cap{T}_{r}|/|{T}_{r}|)$ and Klein et al. 5 define it as $(\sum_{r}|{S}_{r}\cap{T}_{r}|)/(\sum_{r}|{T}_{r}|)$, likewise for other metrics. The latter aggregation is denoted with the suffix (Klein) in the figure. In all four datasets, the boxplots show a narrower interquartile range and substantially higher median than ANTs (higher is better), underscoring the stability and accuracy of our algorithm. (b): Other measures of anatomical label overlap used in 5 are false positives ($|{T}_{r}\backslash{S}_{r}|/|{T}_{r}|$), false negatives ($|{S}_{r}\backslash{T}_{r}|/|{S}_{r}|$), and volume similarity ($2(|{S}_{r}|-|{T}_{r}|)/(|{S}_{r}|+|{T}_{r}|)$) (lower is better). We observe similar trends as in (a), with a narrower interquartile range and substantially lower median values. Results of per region overlap metrics are in the  Fig. S.1.
@@
(a) Trick to avoid parallel transport in Riemannian Adaptive Optimization
@@
Figure S.3: Effect of downsampling on the warp and determinant of the Jacobian: We show the effect of downsampling on the warp and determinant of the Jacobian for a single image pair. The first column shows the initial warp, and the second and third columns show the warp and determinant of the Jacobian for the cubic and bilinear interpolations, respectively.
Figure S.4: Qualitative results on EMPIRE10 challenge: (a) shows the fixed image, (b) shows the registration performed by ANTs, and (c) our method, all with zoomed in regions. ANTs performs a coarse registration with ease, but still leaves out critical alignment of lung boundary and airways by not utilizing adaptive optimization. Our method performs perfectly diffeomorphic registration by construction, and does not lead to any registration errors, both in the lung boundaries or internal features.
Figure S.5: More Qualitative results on EMPIRE10 challenge: (a) shows the fixed image, (b) shows the registration performed by ANTs, and (c) our method, all with zoomed in regions. ANTs performs a coarse registration with ease, but still leaves out critical alignment of lung boundary and airways by not utilizing adaptive optimization. Our method performs perfectly diffeomorphic registration by construction, and does not lead to any registration errors, both in the lung boundaries or internal features.
Figure S.6: Comparison of exponential versus direct optimization on LPBA40 dataset: We run the hyperparameter grid search on the LPBA40 dataset using direct Riemannian gradient updates with Adam optimizer (denoted as rgd), and optimizing the velocity field by computing the exponential map to represent the diffeomorphism (denoted as exp) across all the configurations shown in Fig. 6(a). The average target overlap for each configuration is then stored, and a histogram of target overlap values of the dataset is constructed. Note that the rgd variant has a significantly more number of configurations near the optimal value, and the average performance and the overall distribution of our optimization is better for the rgd variant than exp. Similar trends can be observed for the EMPIRE10 lung challenge in Fig. 3, where the exp representation underperforms for the same cost function, data, etc. Therefore, we recommend direct RGD optimization for diffeomorphisms.

### v2

@@ 4.5 Interpolation strategies for multi-scale registration
$\varphi_{k}=\begin{cases}{id}&\text{if }k=1\\ \text{Upsample}(\varphi_{k-1})&\text{otherwise}\end{cases}$
@@
Unlike upsampling images, upsampling warp fields and their corresponding optimizer state variables requires careful consideration of the interpolation strategy. Bicubic interpolation is a commonly used strategy for upsampling images to preserve smoothness and avoid aliasing. However, bicubic interpolation of the warp field can lead to overshooting, leading to introducing singularities in the upsampled displacement field when there existed none in the original displacement field. In contrast, bilinear or trilinear interpolation does not lead to overshooting, and therefore diffeomorphism of the upsampled displacement is guaranteed, if the original displacement is diffeomorphic. We demonstrate this using a simple 2D warp field in  Fig. 12(b). On the left, we consider a warp field created by nonlinear shear forces. This warp field does not contain any tears or folds - and is diffeomorphic. We upsample this warp field using bicubic interpolation (top) and bilinear interpolation (bottom). We also plot a heatmap of the negative of the determinant of the Jacobian of the upsampled warp, with a white contour representing the zero level set. Qualitatively, bicubic interpolation introduces noticable folds in the warping field, leading to non-diffeomorphisms in the upsampled warp field. The heatmap shows a significant portion of the upsampled warp field has a negative determinant, indicating non-invertibility. On the other hand, bilinear interpolation looks blocky but preserves diffeomorphism everywhere, as also quantitatively verified by the absence of a zero level set in the heatmap.
The complete algorithm is described in  Algorithm 1.
Algorithm 1 FireANTs
1:Input: Fixed image $I_{f}$, Moving image $I_{m}$
2:Scales $[s_{1},s_{2},\ldots,s_{n}]$, Iterations $[T_{1},T_{2},\ldots T_{n}]$, $n$ scales
3:optstate optimizer state (for Adam, RMSProp, etc.)
4:use_jac boolean specifying whether to use Jacobian in descent direction
5:
6:Initialize $\varphi\leftarrow\mathbf{id}_{s_{1}}$. $\triangleright$ Initialize warp to identity at first scale
7:Initialize ${l}\leftarrow 1$. $\triangleright$ Initialize current scale
8:while ${l}\leq n$ do
9:     Initialize $i\leftarrow 0$
10:     Initialize $I_{f}^{l},I_{m}^{l}\leftarrow\text{downsample}(I_{f},s_{l}),\text{downsample}(I_{m},s_{l})$
11:     while $i<T_{{l}}$ do
12:         $U_{i}\leftarrow C(I_{f}^{l},I_{m}^{l}\circ\varphi^{i})+R(\varphi)$
13:         Compute $v_{d}^{\prime}(x)\leftarrow\frac{\partial U_{i}}{\partial\varphi}(\varphi^{(i)})(x)$triangleright Jacobian-free Eulerian descent direction
14:         if use_jac then
15:              Compute $v_{d}^{\prime}(x)\leftarrow J^{T}(\varphi^{(i)}(x))v_{d}^{\prime}(x)$triangleright Eulerian descent direction
16:         end if
17:         Update $(v_{d}^{\prime}(x),\text{{optstate}})\leftarrow\text{{optstate}}(v_{d}^{\prime}(x))$triangleright Apply and update optimizer state
18:         Update $\varphi^{(i+1)}\leftarrow\varphi^{(i)}\circ\exp_{id}(\epsilon_{i}v_{d}^{\prime})\approx\varphi^{(i)}\circ(id+\epsilon_{i}v_{d}^{\prime})$
19:         $i\leftarrow i+1$
20:     end while
21:     if $l<n$ then
22:         $\varphi\leftarrow\text{Upsample}(\varphi,s_{(l+1)})$triangleright Upsample warp to next scale using bilinear/trilinear interpolation
23:     end if
24:     $l\leftarrow l+1$
25:end while
Figure 7: Algorithm for FireANTs Algorithm 1 outlines the key steps in FireANTs - computing the Jacobian-free Eulerian descent direction which is simply the Gateaux derivative. If the boolean use_jac is specified, then use the steepest Eulerian descent direction instead. This descent direction is then modified using any adaptive optimization algorithm denoted as optstate. The warp field is then updated using the exponential map or retraction map for small $\epsilon_{i}$. After optimization at a given scale, the warp field is upsampled using bilinear or trilinear interpolation to the next scale until optimization is complete for all steps.
Appendix A Supplementary Information
A.1 Datasets used in Klein et al.challenge
Brain image data and their corresponding labels for 80 normal subjects were acquired from four different datasets. The LPBA40 dataset contains 40 brain images and their labels to construct the LONI Probabilistic Brain Atlas (LPBA40). All volumes were skull-stripped, and aligned to the MNI305 atlas  70 using rigid-body transformation to correct for head tilt. For all these subjects, 56 structures were manually labelled and bias-corrected using the BrainSuite software. The IBSR18 dataset contains brain images acquired at different laboraties through the Internet Brain Segmentation Repository. The T1-weighted images were rotated to be in Talairach alignment and bias-corrected. Manual labelling is performed resulting in 84 labeled regions. For the CUMC12 dataset, 12 subjects were scanned at Columbia University Medical Center on a 1.5T GE scanner. Images were resliced, rotated, segmented and manually labeled, leading to 128 labeled regions. Finally, the MGH10 dataset contains 10 subjects who were scanned at the MGH/MIT/HMS Athinoula A. Martinos Center using a 3T Siemens scanner. The data is bias-corrected, affine-registered to the MNI152 template, and segmented. Finally the images were manually labeled, leading to 74 labeled regions. All datasets have a volume of $256\times 256\times\{128,124\}$ voxels with varying amounts of anisotropic voxel spacing, ranging from $0.84\times 0.84\times 1.5$mm to $1\times 1\times 1.33$mm.
A.2 Modular software implementation to enable effective experimentation
@@
Our software is implemented completely using default primitives in PyTorch. All code and example scripts is available at
A.3 On the Ill-conditioning of Image Registration
Image registration is a highly ill-conditioned, and non-convex problem necessitating advanced optimization methods for convergence. To provide more intuition on the effect of $\kappa$ on convergence of the SGD algorithm, we consider a toy example of a 2D optimization problem. Specfically, we consider a loss function $f_{\kappa}(x,y)=x^{2}+\kappa y^{2}$ where $\kappa>1$ becomes the condition number of the problem. Qualitatively, the effect of the first term diminishes exponentially fast with $\kappa$ (Fig. 8(a)). Quantitatively, we run both SGD and Adam optimization for a 1000 iterations starting from the point $(x,y)=(5,5)$. Fig. 8(c) shows that SGD works extremely well for $\kappa=1$ which is the best-conditioned loss function, but quickly gets stuck for $\kappa\geq 100$. On the contrary, Adam is invariant to the condition number and converges to the minima for all values of $\kappa$. This is because for a diagonal Hessian (as in this case), the second-order adaptive terms are proportional to the diagonal elements of the Hessian. These condition numbers are vanishingly small compared to those in typical image registration tasks, which can exceed $10^{5}$, making them extremely ill-conditioned.
We also consider a more realistic, but tractable scenario of the convex loss function $f_{\kappa,\theta}(x,y)=x_{\theta}^{2}+\kappa y_{\theta}^{2}$, where
$\begin{bmatrix}x_{\theta}\\ y_{\theta}\end{bmatrix}=\begin{bmatrix}\cos(\theta)&\sin(\theta)\\ -\sin(\theta)&\cos(\theta)\end{bmatrix}\begin{bmatrix}x\\ y\end{bmatrix}$
We choose $\theta=\pi/3$ for this experiment. This is simply a rotated version of the previous family of loss functions, as shown in Fig. 8(b). The trajectories obtained from optimization using SGD (Fig. 8(d)) are virtually identical to that in Fig. 8(c) since the new gradients are simply rotated versions of the previous gradients, and the distance from the minima is invariant to the rotation. However, the trajectories from Adam optimization are qualitatively very different, owing to the increasing difference between the true Hessian and its diagonal approximation. Even so, the final point is at a distance of less than $10^{-3}$ units to the minima for $\kappa=1000$, showing the effectiveness of adaptive optimization even for ill-conditioned, non-diagonal Hessians. This is a strong motivation to extend adaptive optimization for non-Euclidean diffeomorphic registration, which is very high-dimensional and ill-conditioned.
(a) Log-Loss landscape of the toy problem $f_{\kappa}(x,y)=x^{2}+\kappa y^{2}$ for $\kappa=1,10,100,1000$. The log-loss becomes increasingly sharp along the y-direction as $\kappa$ increases.
(b) Log-Loss landscape of the toy problem $f_{\kappa}(x,y)=x_{\theta}^{2}+\kappa y_{\theta}^{2}$ for $\kappa=1,10,100,1000$, where $(x_{\theta},y_{\theta})$ is the coordinate $(x,y)$ rotated by an angle $\theta$ about the origin.
(c) Optimization of Fig. 8(a) using SGD and Adam shows that SGD fails to recover the minima for $\kappa\geq 100$ while Adam is invariant to the condition number for diagonal Hessian matrices. This is a strong motivation to use first order adaptive optimization for registration where the condition number can exceed $10^{5}$.
(d) Optimization of Fig. 8(b) using SGD shows identical optimization trajectories as Fig. 8(c). Adam, however, is not invariant to the condition number because the difference between the true Hessian and its diagonal approximation increases with $\kappa$. Even so, the final point is at a distance of less than $10^{-3}$ units to the minima, showing the mitigating effect of adaptive optimization even for non-diagonal Hessians.
Figure 9: Overview of FireANTs and its features: (a) shows the modalities our method is tested on. We demonstrate results on in-vivo T1-weighted brain MRI, lung CT, and expansion microscopy volumes. FireANTs can optimize intensity images as well as binary masks (lung masks in CT) or entire anatomical label maps (brain MRI). (b) shows the technical contributions of FireANTs. We extend Adaptive Optimization to multi-scale Diffeomorphisms by first writing the Riemannian gradient update, and then avoiding parallel transport of the optimization state by leverging the interchangability of the Riemannian gradient at arbitrary transform $\varphi_{t}$ with the Riemannian gradient at $\varphi=\textbf{Id}$. For the Lie-algebra representation, the Gateaux derivative $\frac{\partial L}{\partial\varphi}$ is projected to $\frac{\partial L}{\partial v}$ using analytical backprop. Since the Lie algebra is a vector space, we use standard adaptive optimizers (see Section 4.3.2 for more details). (c) takes a closer look at multi-scale interpolation for diffeomorphisms represented as a warp field. Bicubic interpolation can introduce folding of the warp field at a finer resolution due to overshooting, but bilinear interpolation does not. Therefore, we use this for interpolating the warp field and the optimizer state. (d) shows the extensive experimental setup. Our method is orders of magnitude faster, has state-of-the-art performance on 3 challenges, is robust across hyperparameters and datasets, and is modular and easy to extend.
@@ A.3 On the Ill-conditioning of Image Registration
Figure 10: Regionwise target overlap on the brain MRI datasets: We further evaluate regionwise overlap scores by sampling 15 regions from each dataset, and comparing their distribution using our method and ANTs. Our method has a much higher median score, and better interquartile ranges across regions, demonstrating both accuracy and robustness.
@@
Figure 11: Comparison of our method with ANTs on 4 MRI brain datasets: Registration quality is validated by measuring volume overlap of label maps between the fixed and warped label maps. (a): For anatomical region $r$, warped (binary) label map ${S}_{r}$ and fixed label map ${T}_{r}$, target and mean overlap are defined as $|{S}_{r}\cap{T}_{r}|/|{T}_{r}|$ and $2|{S}_{r}\cap{T}_{r}|/(|{S}_{r}|+|{T}_{r}|)$. We define the aggregate target overlap over all anatomical regions as $\sum_{r}(|{S}_{r}\cap{T}_{r}|/|{T}_{r}|)$ and Klein et al. 29 define it as $(\sum_{r}|{S}_{r}\cap{T}_{r}|)/(\sum_{r}|{T}_{r}|)$, likewise for other metrics. The latter aggregation is denoted with the suffix (Klein) in the figure. In all four datasets, the boxplots show a narrower interquartile range and substantially higher median than ANTs (higher is better), underscoring the stability and accuracy of our algorithm. (b): Other measures of anatomical label overlap used in 29 are false positives ($|{T}_{r}\backslash{S}_{r}|/|{T}_{r}|$), false negatives ($|{S}_{r}\backslash{T}_{r}|/|{S}_{r}|$), and volume similarity ($2(|{S}_{r}|-|{T}_{r}|)/(|{S}_{r}|+|{T}_{r}|)$) (lower is better). We observe similar trends as in (a), with a narrower interquartile range and substantially lower median values. Results of per region overlap metrics are in the  Fig. 10.
@@
(a) Trick to avoid parallel transport in Riemannian Adaptive Optimization using Eulerian differentials
@@
Figure 12: Overview of tricks for multi-scale adaptive optimization for diffeomorphisms: (a) We exploit the group structure of diffeomorphisms to define an Eulerian differential that avoids the need for parallel transport in adaptive optimization algorithms. (b) We show the effect of downsampling on the warp and determinant of the Jacobian for a single image pair. The first column shows the initial warp, and the second and third columns show the warp and determinant of the Jacobian for the cubic and bilinear interpolations, respectively.
Figure 13: Qualitative results on EMPIRE10 challenge: (a) shows the fixed image, (b) shows the registration performed by ANTs, and (c) our method, all with zoomed in regions. ANTs performs a coarse registration with ease, but still leaves out critical alignment of lung boundary and airways by not utilizing adaptive optimization. Our method performs perfectly diffeomorphic registration by construction, and does not lead to any registration errors, both in the lung boundaries or internal features.
Figure 14: More Qualitative results on EMPIRE10 challenge: (a) shows the fixed image, (b) shows the registration performed by ANTs, and (c) our method, all with zoomed in regions. ANTs performs a coarse registration with ease, but still leaves out critical alignment of lung boundary and airways by not utilizing adaptive optimization. Our method performs perfectly diffeomorphic registration by construction, and does not lead to any registration errors, both in the lung boundaries or internal features.
Figure 15: Comparison of exponential versus direct optimization on LPBA40 dataset: We run the hyperparameter grid search on the LPBA40 dataset using direct Riemannian gradient updates with Adam optimizer (denoted as rgd), and optimizing the velocity field by computing the exponential map to represent the diffeomorphism (denoted as exp) across all the configurations shown in Fig. 5(a). The average target overlap for each configuration is then stored, and a histogram of target overlap values of the dataset is constructed. Note that the rgd variant has a significantly more number of configurations near the optimal value, and the average performance and the overall distribution of our optimization is better for the rgd variant than exp. Similar trends can be observed for the EMPIRE10 lung challenge in Fig. 2, where the exp representation underperforms for the same cost function, data, etc. Therefore, we recommend direct RGD optimization for diffeomorphisms.